The Florida Investment Fund buys 58 bonds of the Gator Corporation through a broker. The bonds pay 10 percent annual interest. The yield to maturity (market rate of interest) is 12 percent. The bonds have a 10-year maturity. Using an assumption of semiannual interest payments: Compute the price of a bond (refer to “Semiannual Interest and Bond Prices” in Chapter 10 for review if necessary). Compute the total value of the 58 bonds.
Answer #1

When the interest is paid half yearly, then the interest rates are reduced to half and period gets double.

The price of the bond = half year's interest * PVIFA(Yield interest, n) + Principle * PVIF(Yield interest,n)
Half year interest = 1000 * (10% * 1/2) = 50
PVIFA(Yield interest, n) = PVIFA(6%, 20)= 11.4699
PVIF(Yield interest,n) = PVIF(6%,20) = 0.3118
Principle = $1000
The price of the bond = 50 * 11.4699 + 1000 * 0.3118 = 885.3
The total value of the 58 bonds = 885.30 * 58 = $ 51347.40